Do flexible schedules reduce the demand for​ resources? A government health department experiments with a flexible​ four-day work week. For a​ year, the department recorded the mileage driven by 11 field workers on an ordinary​ five-day work week.​ Then, it changed to a flexible​ four-day work week and recorded mileage for another year. The resulting sample had a standard deviation of 1111.3. How big should the sample have been to narrow the width of the reported​ 95% confidence interval to 600 miles? Use the sample standard deviation as your standard deviation guess.
Answer #1

Solution:

Given:

Standard deviation = s =  1111.3

c = confidence level = 95%

Width of confidence interval = 600 miles

then

Margin of Error = E= Width / 2 = 600 / 2= 300 miles

We have to find sample size n.

Formula:

n= \left ( \frac{Z_{c}\times \sigma }{E} \right )^{2}

\sigma is unknown , but we use its estimate sample standard deviation s.

We need to find zc value for c=95% confidence level.

Find Area = ( 1 + c ) / 2 = ( 1 + 0.95) /2 = 1.95 / 2 = 0.9750

Look in z table for Area = 0.9750 or its closest area and find z value.

0.1 V123 .00 .01 .02 .03 .04 .05 .06 0.0 ,5000 15040 5080 5120 .5160 5199 39 1.5398 .5438 5478 1.5517 .5557 .5596 .636 0.2 15

Area = 0.9750 corresponds to 1.9 and 0.06 , thus z critical value = 1.96

That is : Zc = 1.96

Thus

n=(2x) n = (Zc X 5 Ε

(1.96 x 1111.3 300 )

n = (7.26049)

n = 52.71476

n = 53